Find the order and degree of the following differential equation 2(d^2y/dx^2)-3*(dy/dx)+y=0


1
Expert's answer
2022-03-30T17:05:17-0400

"2(\frac{d^2y}{dx^2})-3(\frac{dy}{dx})+y=0"

Highest order derivative="\frac{d^2y}{dx^2}"

So order of the differential equation = 2


The highest order derivative is raised to power 1

So, the degree of the differential equation = 1